# Physics-informed Value Learner for Offline Goal-Conditioned Reinforcement Learning

### v1

@@ 4 Physics-informed Eikonal regularizer
As discussed in Remark 4.3, the effectiveness of the regularization in Eq. (9) stems from the geometric structure of the Eikonal PDE, whose solution defines a signed distance field Ni and Qureshi (2023); Smith et al. (2020). In our final formulation, we uniformly set the speed profile to $S(s)=1$, corresponding to the constraint $\|\nabla_{s}V_{\bm{\theta}_{V}}(s,g)\|\approx 1$ across the feasible state space. By Rademacher’s theorem Heinonen (2005), this enforces 1-Lipschitz continuity almost everywhere, encouraging smoother and more generalizable value estimates even under limited offline data coverage. Although the true optimal value function $V^{*}$ may not be exactly 1-Lipschitz over the entire feasible space, we show in our experiments (Section 5) that this regularization remains effective in practice. Note that using a different constant $S(s)>0$ everywhere in the feasible space would simply uniformly rescale the value function without affecting its shape. Consequently, since policy gradients are invariant to such rescaling Sutton et al. (1999b); Schulman et al. (2015); Sutton and Barto (2018), this design choice has no impact on the induced policy.
Physics-informed hierarchical implicit Q-learning
@@
In the following, we provide a brief summary of our algorithm, Pi-HIQL. During training, Pi-HIQL performs a Pi value estimation step, where the loss in (8)-(9) is minimized to learn a GCVF. This is followed by a policy extraction step, in which the hierarchical actor introduced in Park et al. (2024b) (see Preliminaries in Section 3) is trained based on the estimated GCVF. The full pseudo-code for Pi-HIQL as well as a JAX Bradbury et al. (2018) implementation showing how to compute the gradient $\nabla_{s}V_{\bm{\theta}_{V}}$ in (9) are provided in Appendix C, respectively Algorithm 1 and Algorithm 2. For the full implementation of Pi-HIQL refer to our GitHub repository.
@@ 5 Experiments
Figure 2: Environments from OGbench Park et al. (2024c) used in our experiments. These include a variety of goal-conditioned tasks spanning navigation and locomotion (e.g., pointmaze, antmaze, humanoidmaze), contact-rich locomotion (antsoccer), and contact-rich manipulation (cube, scene). The environments differ significantly in dynamics complexity, dimensionality, and task structure, providing a comprehensive testbed for evaluating Offline GCRL algorithms.
In this section, we analyze the effects of the Eikonal regularizer in (9) on the GCVF estimation problem. Specifically, we will perform an ablation over different designs of speed profiles $S(s)$, compare the Eikonal regularizer with an HJB regularizer, and analyze value functions learned with and without our Pi term. Then, we compare the performance obtained by our Eikonal-regularized algorithm, Pi-HIQL, against the SOTA algorithms for Offline GCRL. Finally, we will also present the limitations of our approach. The experiments in this section are conducted on the environments in Fig. 2. A table summarizing the most relevant hyperparameter values is provided in Appendix C.
@@
Table 1: Summary of the speed profiles ablation. All agents are trained for 100,000 training steps using 10 seeds. We report the mean and standard deviation across seeds for the best evaluation achieved during training. For each seed, evaluations are conducted over $5$ different random goals, as designed in Park et al. (2024c), with the learned policy tested for $50$ episodes per goal. Results within $95\%$ of the best value are written in bold.
Environment Dataset Type Maze Dimension Pi-HIQL Pi-HIQL Exp (10) Pi-HIQL Lin (11) Pi-HIQL HJB (12)
@@
Recall that our algorithm, Pi-HIQL, estimates the GCVF using the Pi loss defined in Eqs. (8)-(9), where the speed profile $S(s)$ encodes task-specific structure into the Eikonal regularizer. We perform an ablation over different choices of $S(s)$ and find the simple constant setting $S(s)=1$ to be particularly effective. It outperforms more complex alternatives that require explicit knowledge of obstacle coordinates, which may not be available in practice. Specifically, we compare $S(s)=1$ against the following two speed profiles:
@@
where $\lambda$ is a decay rate, $S_{\text{min}}$ represents the minimum tolerable speed, $d_{\text{min}}$ and $d_{\text{max}}$ are respectively the minimum and maximum tolerable distances of a state $s$ from its nearest obstacle and $d(s)$ is a function $d:\mathcal{S}\to\mathbb{R}$ describing the Euclidean distance of the state $s$ from its nearest obstacle (see Appendix B for more information). We refer to $S_{\text{exp}}(\cdot)$ in (10) as the exponential speed profile (Pi-HIQL Exp in Table 1) and to $S_{\text{lin}}(\cdot)$ in (11), originally introduced in Ni and Qureshi (2023), as the linear speed profile (Pi-HIQL Lin in Table 1). Both (10) and (11) assign high speed values to states $s$ far from obstacles, and low speed values to states near obstacles. This encourages the agent to avoid obstacles, as proximity to them results in longer travel-time from $s$ to $g$. Moreover, due to the use of the $\exp(\cdot)$ function, when compared to $S_{\text{lin}}(\cdot)$, $S_{\text{exp}}(\cdot)$ ensures a smoother decay of the speed value as the distance from the obstacles decreases Ni and Qureshi (2024). Finally, note that, in both these speed profiles, computing $d(s)$ requires knowledge of the obstacles’ coordinates which represents a strong assumption in some settings. In addition to this ablation, we also compare our Eikonal regularization term in (9) with an HJB regularizer derived from the HJB PDE in (5):
@@
In (12), the dynamics $f(s,a)$, originally required by the HJB PDE in (5), is replaced by a finite difference term $(s^{\prime}-s)$ as proposed by Lien et al. (2024). The term in (12) is then used in place of the Eikonal regularizer in (9) for GCVF estimation in Pi-HIQL. We refer to this approach as Pi-HIQL HJB in Table 1. The results for these experiments are summarized in Table 1 and all the learning curves are available in Appendix D.
Two main observations explain the superior performance of the simple choice $S(s)=1$. First, in Offline GCRL, sampled trajectories are already constrained to the feasible space, and there is no explicit penalty for colliding with obstacles during learning. As such, complex speed profiles incorporating obstacle proximity provide limited additional benefit over the uniform case. Second, the simplicity of $S(s)=1$ enables more efficient and stable learning of the GCVF, whereas profiles requiring privileged information (e.g., obstacle maps) may introduce unnecessary complexity. Our choice of $S(s)=1$ also aligns with standard practice in the literature, which commonly adopts uniform speed profiles to model wavefront propagation in the feasible state space Sethian (1996); Ni and Qureshi (2023). This also ensures a fair comparison with the Offline GCRL baselines, as none of them leverage privileged information. Furthermore, we provide visualizations for the learned GCVF in Appendix B, where contour plots for pointmaze-giant-stitch-v0 illustrate that Pi-HIQL with $S(s)=1$ learns a more structured and accurate value function, closely aligning with the maze layout. In contrast, Pi-HIQL Exp and Pi-HIQL Lin display artifacts even near the goal, while Pi-HIQL HJB fails to recover the maze geometry altogether.
Based on this empirical evidence, and to ensure a fair comparison with baselines, we adopt the constant speed profile $S(s)=1$ in all subsequent experiments. Nonetheless, we highlight that in the context of Safe GCRL—where value functions must encode both safety and task performance—investigating how different choices of $S(s)$ influence the shape and behavior of the GCVF represents an interesting direction for future work.
Table 2: Complete comparison between P-HiQRL and the offline GCRL baselines. Agents are trained for 100,000 steps on pointmaze tasks and 1 million steps on the remaining tasks, each using 10 seeds. The evaluation follows the methodology described in Table 1. We report the mean and standard deviation across seeds for the best evaluation achieved during training. Results within $95\%$ of the best value are written in bold, and rows are highlighted when the Eikonal regularizer improves performance by $100\%$ or more compared to the non-regularized HIQL performance.
Environment Dataset Dimension Pi-HIQL HIQL QRL CRL
@@
Pi-HIQL vs HIQL
We compare Pi-HIQL with its non-Pi counterpart, HIQL Park et al. (2024b), to isolate the effect of the Eikonal regularizer. This comparison is conducted under tightly controlled conditions: both methods use identical network architectures, hyperparameters, and training pipelines, ensuring that the only difference lies in the presence of the regularizer. As shown in Table 2, Pi-HIQL consistently outperforms HIQL on navigation tasks, where the value function is typically Lipschitz continuous. The benefits are particularly pronounced in large mazes and stitching regimes, with improvements exceeding 100% in 7 out of 31 evaluated tasks.
Given the magnitude of these gains relative to the standard deviations across 10 random seeds, the improvements are both practically meaningful and statistically significant. Full learning curves for these experiments are provided in Appendix D. These results highlight the robustness of the Eikonal regularizer in smooth environments and demonstrate its ability to enhance generalization in settings where HIQL struggles to scale. As illustrated in Fig. 1 for the antmaze-navigate-giant-v0 task, Pi-HIQL produces a GCVF that reflects the underlying maze geometry, whereas HIQL fails to capture this structure, leading to poor goal-directed performance.
By contrast, in interactive, contact-rich domains such as antsoccer and manipulation, where the dynamics (and consequently the value function) exhibit discontinuities, Pi-HIQL offers limited advantage. We defer a detailed discussion of these cases to the Limitations section below.
@@
We extend our comparison to state-of-the-art Offline GCRL algorithms, including QRL Wang et al. (2023) and CRL Eysenbach et al. (2022), alongside HIQL. Results across all 31 tasks are summarized in Table 2, with full learning curves in Appendix D. Pi-HIQL consistently outperforms all baselines in the most challenging settings—most notably in the antmaze-stitch tasks, which combine complex locomotion with, composite datasets, and in humanoidmaze, where high-dimensional states and unstable control amplify learning difficulty.
QRL performs well in simpler domains such as pointmaze, where its quasimetric structure aids goal-conditioned estimation. CRL is competitive on several navigate variants, but struggles in high-dimensional and stitched tasks such as humanoidmaze. In contrast, Pi-HIQL demonstrates strong generalization and planning performance in both long-horizon and large-scale environments.
In antsoccer and manipulation, however, Pi-HIQL performs comparably to the baselines. As mentioned, these domains involve discontinuous or contact-rich dynamics, and we do not observe consistent benefits from the Eikonal regularizer due to the fact that the imposed gradient condition does not hold globally at optimality throughout the entire state space.
(a) Pi-HIQL(b) HIQL
Figure 3: Countour plots of the GCVF on antsoccer-medium-navigate-v0 Park et al. (2024c), learned after 1 million steps by Pi-HIQL and HIQL respectively. These plots are generated following the same methodology in Fig. 1.
@@
We conclude by discussing some limitations of our approach, particularly in contact-rich tasks that involve interactions with external objects. In the antsoccer domain, for instance, an Ant agent must not only navigate but also interact with a soccer ball to reach a specified goal. Similarly, in manipulation tasks such as cube-single-play, the agent must coordinate precise contact interactions with objects in the scene. These tasks introduce hybrid, non-Lipschitz dynamics due to discontinuities in contact states—often modeled as categorical variables (e.g., in-contact vs. free motion)—which pose a challenge for regularizers grounded in smoothness and Lipschitz continuity assumptions Kim et al. (2021), such as our Eikonal term. As visualized in Fig. 3 for the antsoccer experiments, Pi-HIQL learns a GCVF that aligns well with the environment geometry, particularly for the navigational components. However, this structured value function does not consistently yield better performance (cf. Table 2) as these tasks also require complex interactions with objects in the environment. Similar trends are observed in manipulation tasks, where Pi-HIQL performs comparably to the baselines but does not show marked gains. These experiments are included for completeness and highlight that, while Pi-HIQL excels in navigation-dominated domains, additional mechanisms—such as task-adaptive speed profiles or representation learning tailored to contact dynamics—may be necessary to extend its benefits to interactive, contact-rich environments. We leave this direction for future work.
@@ 6 Conclusion
We introduced Pi-HIQL, a novel approach to Offline GCRL that integrates an Eikonal PDE-based regularizer for GCVF estimation. Our analysis demonstrated that the Eikonal regularizer effectively introduces a useful distance-like inductive bias, which promotes consistent gradient magnitudes and improves value estimation in high-dimensional spaces (cf. Fig. 1). This property mitigates irregularities from the limited coverage of offline datasets, resulting in robust, globally consistent GCVFs that accurately capture the underlying structure of the environment. Consequently, Pi-HIQL outperformed SOTA baselines across diverse tasks, excelling in complex scenarios such as large-scale mazes and trajectory stitching, where traditional methods often fail to generalize.
@@ Appendix B Speed profiles and HJB comparison: additional details and contour plots
Figure 4: Fig. 4(a) illustrates how to compute the distance function $d(s)$ used in (10) and (11). Given $s$, we first determine the nearest obstacle in terms of the euclidean distance and then set $d(s)$ equal to this distance. Fig. 4(b) shows speed results obtained using $S_{\text{lin}}(s)$ in (11) for the pointmaze-medium-navigate-v0 dataset.
(a) Pi-HIQL(b) Pi-HIQL Exp (10)(c) Pi-HIQL Lin (11)(d) Pi-HIQL HJB (12)
Figure 5: Contour plots of the GCVF on pointmaze-giant-stitch-v0 Park et al. [2024c], learned after 100,000 training steps by the algorithms in Table 1. We notice that Pi-HIQL, with constant speed profile $S(s)=1$, provides the best GCVF estimation which results in the highest score for the specific task.
Appendix C Psuedocode and Hyperparameters
During training, Pi-HIQL performs a Pi value estimation step followed by a hierarchical policy extraction step. During the value estimation step, the following loss, as provided in (8)-(9), is minimized:

### v3

@@ 4 Physics-informed Eikonal regularizer
As discussed in Remark 4.3, the effectiveness of the regularization in Eq. (9) stems from the geometric structure of the Eikonal PDE, whose solution defines a signed distance field [37, 36]. In our final formulation, we uniformly set the speed profile to $S(s)=1$, corresponding to the constraint $\|\nabla_{s}V_{\bm{\theta}_{V}}(s,g)\|\approx 1$ across the feasible state space. By Rademacher’s theorem [46], this enforces 1-Lipschitz continuity almost everywhere, encouraging smoother and more generalizable value estimates even under limited offline data coverage. Although the true optimal value function $V^{*}$ may not be exactly 1-Lipschitz over the entire feasible space, we show in our experiments (Section 5) that this regularization remains effective in practice. Note that using a different constant $S(s)>0$ everywhere in the feasible space would simply uniformly rescale the value function without affecting its shape. Consequently, since policy gradients are invariant to such rescaling [47, 48, 42], this design choice has no impact on the induced policy.
Eikonal-regularized HIQL
@@
In the following, we provide a brief summary of our algorithm, Eik-HIQL. During training, Eik-HIQL performs an Eikonal-regularized value estimation step, where the loss in (8)-(9) is minimized to learn a GCVF. This is followed by a policy extraction step, in which the hierarchical actor introduced in Park et al. [10] (see Preliminaries in Section 3) is trained based on the estimated GCVF. The full pseudo-code for Eik-HIQL as well as a JAX [49] implementation showing how to compute the gradient $\nabla_{s}V_{\bm{\theta}_{V}}$ in (9) are provided in Appendix D, respectively Algorithm 1 and Algorithm 2. For the full implementation of Eik-HIQL refer to our GitHub repository.
@@ 5 Experiments
Figure 2: Environments from OGbench [11] used in our experiments. These include a variety of goal-conditioned tasks spanning navigation and locomotion (e.g., pointmaze, antmaze, humanoidmaze), contact-rich locomotion (antsoccer), and contact-rich manipulation (cube, scene). The environments differ significantly in dynamics complexity, dimensionality, and task structure, providing a comprehensive testbed for evaluating Offline GCRL algorithms.
In this section, we analyze the effects of the Eikonal regularizer in (9) on the GCVF estimation problem. Specifically, we will perform an ablation over different designs of speed profiles $S(s)$, compare the Eikonal regularizer with an HJB regularizer, and analyze value functions learned with and without our Eikonal term. Then, we compare the performance obtained by our Eikonal-regularized algorithm, Eik-HIQL, against the SOTA algorithms for Offline GCRL. Finally, we will also present the limitations of our approach. The experiments in this section are conducted on the environments in Fig. 2. A table summarizing the most relevant hyperparameter values is provided in Appendix D.
@@
Table 1: Summary of the speed profiles ablation. All agents are trained for 100,000 training steps using 10 seeds. We report the mean and standard deviation across seeds for the best evaluation achieved during training. For each seed, evaluations are conducted over $5$ different random goals, as designed in Park et al. [11], with the learned policy tested for $50$ episodes per goal. Results within $95\%$ of the best value are written in bold.
Environment Dataset Type Maze Dimension Eik-HIQL Eik-HIQL Exp (10) Eik-HIQL Lin (11) HJB-HIQL (12)
@@
Recall that our algorithm, Eik-HIQL, estimates the GCVF using the loss defined in Eqs. (8)-(9), where the speed profile $S(s)$ encodes task-specific structure into the Eikonal regularizer. We perform an ablation over different choices of $S(s)$ and find the simple constant setting $S(s)=1$ to be particularly effective. It outperforms more complex alternatives that require explicit knowledge of obstacle coordinates, which may not be available in practice. Specifically, we compare $S(s)=1$ against the following two speed profiles:
@@
where $\lambda$ is a decay rate, $S_{\text{min}}$ represents the minimum tolerable speed, $d_{\text{min}}$ and $d_{\text{max}}$ are respectively the minimum and maximum tolerable distances of a state $s$ from its nearest obstacle and $d(s)$ is a function $d:\mathcal{S}\to\mathbb{R}$ describing the Euclidean distance of the state $s$ from its nearest obstacle (see Appendix B for more information). We refer to $S_{\text{exp}}(\cdot)$ in (10) as the exponential speed profile (Eik-HIQL Exp in Table 1) and to $S_{\text{lin}}(\cdot)$ in (11), originally introduced in [37], as the linear speed profile (Eik-HIQL Lin in Table 1). Both (10) and (11) assign high speed values to states $s$ far from obstacles, and low speed values to states near obstacles. This encourages the agent to avoid obstacles, as proximity to them results in longer travel-time from $s$ to $g$. Moreover, due to the use of the $\exp(\cdot)$ function, when compared to $S_{\text{lin}}(\cdot)$; $S_{\text{exp}}(\cdot)$ ensures a smoother decay of the speed value as the distance from the obstacles decreases [50]. Finally, note that, in both these speed profiles, computing $d(s)$ requires knowledge of the obstacles’ coordinates which represents a strong assumption in some settings. In addition to this ablation, we also compare our Eikonal regularization term in (9) with an HJB regularizer derived from the HJB PDE in (5):
@@
In (12), the dynamics $f(s,a)$, originally required by the HJB PDE in (5), is replaced by a finite difference term $(s^{\prime}-s)$ as proposed by Lien et al. [6]. The term in (12) is then used in place of the Eikonal regularizer in (9) for GCVF estimation in Eik-HIQL. We refer to this approach as HJB-HIQL in Table 1. The results for these experiments are summarized in Table 1 and all the learning curves are available in Appendix E.
Two main observations explain the superior performance of the simple choice $S(s)=1$. First, in Offline GCRL, sampled trajectories are already constrained to the feasible space, and there is no explicit penalty for colliding with obstacles during learning. As such, complex speed profiles incorporating obstacle proximity provide limited additional benefit over the uniform case. Second, the simplicity of $S(s)=1$ enables more efficient and stable learning of the GCVF, whereas profiles requiring privileged information (e.g., obstacle maps) may introduce unnecessary complexity. Our choice of $S(s)=1$ also aligns with standard practice in the literature, which commonly adopts uniform speed profiles to model wavefront propagation in the feasible state space [51, 37]. This also ensures a fair comparison with the Offline GCRL baselines, as none of them leverage privileged information. Furthermore, we provide visualizations for the learned GCVF in Appendix B, where contour plots for pointmaze-giant-stitch-v0 illustrate that Eik-HIQL with $S(s)=1$ learns a more structured and accurate value function, closely aligning with the maze layout. In contrast, Eik-HIQL Exp and Eik-HIQL Lin display artifacts even near the goal, while Eik-HIQL HJB fails to recover the maze geometry altogether. Based on this empirical evidence, and to ensure a fair comparison with baselines, we adopt the constant speed profile $S(s)=1$ in all subsequent experiments. Nonetheless, we highlight that in the context of Safe GCRL, where value functions must encode both safety and task performance, investigating how different choices of $S(s)$ influence the shape and behavior of the GCVF represents an interesting direction for future work.
Table 2: Complete comparison between Eik-HiQRL and the Offline GCRL baselines. Agents are trained for 100,000 steps on pointmaze tasks and 1 million steps on the remaining tasks, each using 10 seeds. The evaluation follows the methodology described in Table 1. We report the mean and standard deviation across seeds for the best evaluation achieved during training. Results within $95\%$ of the best value are written in bold, and rows are highlighted when the Eikonal regularizer improves performance by $100\%$ or more compared to the non-regularized HIQL performance.
Environment Dataset Dim Eik-HIQL HIQL QRL CRL
@@
Eik-HIQL vs HIQL
We compare Eik-HIQL with its non-regularized counterpart, HIQL [10], to isolate the effect of the Eikonal regularizer. This comparison is conducted under tightly controlled conditions: both methods use identical network architectures, hyperparameters, and training pipelines, ensuring that the only difference lies in the presence of the regularizer. As shown in Table 2, Eik-HIQL consistently outperforms HIQL on navigation tasks, where the value function is typically Lipschitz continuous. The benefits are particularly pronounced in large mazes and stitching regimes, with improvements exceeding 100% in 7 out of 31 evaluated tasks.
Given the magnitude of these gains relative to the standard deviations across 10 random seeds, the improvements are both practically meaningful and statistically significant. Full learning curves for these experiments are provided in Appendix E. These results highlight the robustness of the Eikonal regularizer in smooth environments and demonstrate its ability to enhance generalization in settings where HIQL struggles to scale. As illustrated in Fig. 1 for the antmaze-navigate-giant-v0 task, Eik-HIQL produces a GCVF that reflects the underlying maze geometry, whereas HIQL fails to capture this structure, leading to poor goal-directed performance.
By contrast, in interactive, contact-rich domains such as antsoccer and manipulation, where the dynamics (and consequently the value function) exhibit discontinuities, Eik-HIQL offers limited advantage. We defer a detailed discussion of these cases to the Limitations section below.
@@
We extend our comparison to state-of-the-art Offline GCRL algorithms, including QRL [12] and CRL [13], alongside HIQL. Results across all 31 tasks are summarized in Table 2, with full learning curves in Appendix E. Eik-HIQL consistently outperforms all baselines in the most challenging settings, most notably in the antmaze-stitch tasks, which combine complex locomotion with, composite datasets, and in humanoidmaze, where high-dimensional states and unstable control amplify learning difficulty.
QRL performs well in simpler domains such as pointmaze, where its quasimetric structure aids goal-conditioned estimation. CRL is competitive on several navigate variants, but struggles in high-dimensional and stitched tasks such as humanoidmaze. In contrast, Eik-HIQL demonstrates strong generalization and planning performance in both long-horizon and large-scale environments.
In antsoccer and manipulation, however, Eik-HIQL performs comparably to the baselines. As mentioned, these domains involve discontinuous or contact-rich dynamics, and we do not observe consistent benefits from the Eikonal regularizer due to the fact that the imposed gradient condition does not hold globally at optimality throughout the entire state space.
(a) Eik-HIQL(b) HIQL
Figure 3: Countour plots of the GCVF on antsoccer-medium-navigate-v0 [11], learned after 1 million steps by Eik-HIQL and HIQL respectively. These plots are generated following the same methodology in Fig. 1.
@@
We conclude by discussing some limitations of our approach, particularly in contact-rich tasks that involve interactions with external objects. In the antsoccer domain, for instance, an ant agent must not only navigate but also interact with a soccer ball to reach a specified goal. Similarly, in manipulation tasks such as cube-single-play, the agent must coordinate precise contact interactions with objects in the scene. These tasks introduce hybrid, non-Lipschitz dynamics due to discontinuities in contact states, often modeled as categorical variables (e.g., in-contact vs. free motion), which pose a challenge for regularizers grounded in smoothness and Lipschitz continuity assumptions [43], such as our Eikonal term. As visualized in Fig. 3 for the antsoccer experiments, Eik-HIQL learns a GCVF that aligns well with the environment geometry, particularly for the navigational components. However, this structured value function does not consistently yield better performance (cf. Table 2) as these tasks also require complex interactions with objects in the environment. Similar trends are observed in manipulation tasks, where Eik-HIQL performs comparably to the baselines but does not show marked gains. These experiments are included for completeness and highlight that, while Eik-HIQL excels in navigation-dominated domains, additional mechanisms, such as task-adaptive speed profiles or representation learning tailored to contact dynamics, may be necessary to extend its benefits to interactive, contact-rich environments. We leave this direction for future work.
@@ 6 Conclusion
We introduced Eik-HIQL, a novel approach to Offline GCRL that integrates an Eikonal PDE-based regularizer for GCVF estimation. Our analysis demonstrated that the Eikonal regularizer effectively introduces a useful distance-like inductive bias, which promotes consistent gradient magnitudes and improves value estimation in high-dimensional spaces (cf. Fig. 1). This property mitigates irregularities from the limited coverage of offline datasets, resulting in robust, globally consistent GCVFs that accurately capture the underlying structure of the environment. Consequently, Eik-HIQL outperformed SOTA baselines across diverse tasks, excelling in complex scenarios such as large-scale mazes and trajectory stitching, where traditional methods often fail to generalize.
@@ Appendix B Speed profiles and HJB comparison: additional details and contour plots
Figure 4: Fig. 4(a) illustrates the computation of the distance function $d(s)$ used in (10) and (11). Let the state be represented by its spatial coordinates $s=(x,y)\in\mathbb{R}^{2}$, and let $\mathcal{O}=\{o_{1},\ldots,o_{M}\}$ denote the set of obstacle coordinates in the maze. We define $d(s)=\min_{o\in\mathcal{O}}\|s-o\|_{2}$, i.e., the Euclidean distance from $s$ to the nearest obstacle. Fig. 4(b) reports the resulting speed profile obtained using $S_{\text{lin}}(s)$ in (11) for the pointmaze-medium-navigate-v0 dataset.
(a) Eik-HIQL(b) Eik-HIQL Exp (10)(c) Eik-HIQL Lin (11)(d) HJB-HIQL (12)
Figure 5: Contour plots of the GCVF on pointmaze-giant-stitch-v0 [11], learned after 100,000 training steps by the algorithms in Table 1. We observe that Eik-HIQL with constant speed profile $S(s)=1$ provides the most accurate GCVF estimation, resulting in the highest score for this task. In contrast, the contour plot for HJB-HIQL in (d) fails to effectively capture the maze layout. Furthermore, in (b) and (c) we annotate, with red circles, examples of artifacts arising in Eik-HIQL Exp and Eik-HIQL Lin, respectively.
Appendix C HJB PDE step-by-step derivations and Proof of Proposition 4.1
In the following we provide the step-by-step derivations for the HJB PDE in (5) and the full proof for Proposition 4.1.
C.1 HJB PDE derivations
The optimal value function $V(s,g)$ associated with the undiscounted optimal control problem in (3), satisfies the following principle of optimality
where $\Delta t$ is a small time step. By substituting the Taylor expansion
$V(s(t+\Delta t),g)=V(s,g)+\nabla_{s}V(s,g)^{\top}f(s,a)\Delta t+O(\Delta t^{2})$
into (13), we obtain
$V(s,g)=\inf_{a\in\mathcal{A}}[c(s,a)\Delta t+V(s,g)+\nabla_{s}V(s,g)^{\top}f(s,a)\Delta t+O(\Delta t^{2})].$
Subtracting $V(s,g)$ on both sides and dividing by $\Delta t$ gives:
$0=\inf_{a\in\mathcal{A}}[c(s,a)+\nabla_{s}V(s,g)^{\top}f(s,a)+O(\Delta t)].$
Taking the limit $\Delta t\to 0$, we recover Eq. (5):
$\inf_{a\in\mathcal{A}}[c(s,a)+\nabla_{s}V(s,g)^{\top}f(s,a)]=0.$
C.2 Proof of Proposition 4.1
Proof.
Consider the HJB PDE in (5). By applying the Cauchy-Schwarz inequality to the argument of the minimization we obtain
$c(s,a)+\nabla_{s}V(s,g)^{\top}f(s,a)\leq c(s,a)+\|\nabla_{s}V(s,g)\|\|f(s,a)\|.$
Then, by defining
$F^{*}(s)=\sup_{a\in\mathcal{A}}\|f(s,a)\|,$
we can further upper bound the right-hand-side and obtain:
$c(s,a)+\nabla_{s}V(s,g)^{\top}f(s,a)\leq c(s,a)+\|\nabla_{s}V(s,g)\|F^{*}(s).$
Finally, applying the infimum over $a\in\mathcal{A}$ on both sides yields:
$\inf_{a\in\mathcal{A}}[c(s,a)+\nabla_{s}V(s,g)^{\top}f(s,a)]\leq\inf_{a\in\mathcal{A}}[c(s,a)+\|\nabla_{s}V(s,g)\|F^{*}(s)],$
where the result in Eq. (6) is obtained by defining $c^{*}(s)=\inf_{a\in\mathcal{A}}c(s,a)$.
For the equality in (7), note that for the isotropic dynamics $f(s,a)=a$ and given $c(s,a)$ constant over a=1||a||=1, the inner product $\nabla_{s}V(s,g)^{\intercal}f(s,a)$ attains its minimal value when $a$ points in the direction opposite to $\nabla_{s}V(s,g)$. Specifically, this occurs when
[…]
Substituting $f(s,a)=a^{*}$ into
$H(s,g,\nabla_{s}V(s,g))=\inf_{a\in\mathcal{A}}[c(s,a)+\nabla_{s}V(s,g)^{\top}f(s,a)]$
and simplifying yields
[…]
Appendix D Psuedocode and Hyperparameters
During training, Eik-HIQL performs an Eikonal-regularized value estimation step followed by a hierarchical policy extraction step. During the value estimation step, the following loss, as provided in (8)-(9), is minimized: